From the UV absorption data comment on the purity of your DNA preparation.

*Not sure what extra data is required? All data given to us to complete lab is here. 234nm 0.350 260nm 0.524 1st Der. 1st Der. 0.386 0.385 64.02 Temp 55.02 57.02 59.02 61.02 63.02 65.07 67.02 69.04 71.02 73.07Temp 1st Der. Abs gDNA Temp 56.97 57.97 58.97 59.97 60.97 55.22 57.47 59.52 61.52 63.47 65.57 67.57 69.52 71.57 73.52 75.52 7
Answer #1
  • Purity of DNA preparations can be  measured by taking  absorbance from 230nm to 320nm to detect other possible contaminants.
  • However with Absorbance data given here , one can get the idea of DNA purity by applying the most common purity calculation i.e by calculating  ratio of the absorbance at 260nm divided by the reading at 280nm.

Here A260  = 0.524

  A280 = 0.283

DNA purity =  A260/A280

= 0.524/0.283

= 1.851

Good-quality DNA will have ratio  of 1.7–2.0.

In this experiment  it is coming out 1.85 thus Purity of DNA is good.

Part (ii)

Absorbance around 230nm can indicate that organic compounds or chaotropic salts are present in the purified DNA.

Thus ratio of absorbance at 260nm to 230nm gives idea of  the level of salt carryover in the purified DNA.

As absorbance at A234=0.350

In the present experiment A260/A230 = 0.524/0.350 = 1.49

As a guideline, the A260/A230 is best if greater than 1.5 which is slightly lower 1.49 ,

lower ratio, indiate the salt carryover in the purified DNA than  required.